### v1

# Assemblage: Automatic Binary Dataset Construction for Machine Learning

Abstract
Binary code is pervasive, and binary analysis is a key task in reverse engineering, malware classification, and vulnerability discovery. Unfortunately, while there exist large corpuses of malicious binaries, obtaining high-quality corpuses of benign binaries for modern systems has proven challenging (e.g., due to licensing issues). Consequently, machine learning based pipelines for binary analysis utilize either costly commercial corpuses (e.g., VirusTotal) or open-source binaries (e.g., coreutils) available in limited quantities. To address these issues, we present Assemblage: an extensible cloud-based distributed system that crawls, configures, and builds Windows PE binaries to obtain high-quality binary corpuses suitable for training state-of-the-art models in binary analysis. We have run Assemblage on AWS over the past year, producing 890k Windows PE and 428k Linux ELF binaries across 29 configurations. Assemblage is designed to be both reproducible and extensible, enabling users to publish “recipes” for their datasets, and facilitating the extraction of a wide array of features. We evaluated Assemblage by using its data to train modern learning-based pipelines for compiler provenance and binary function similarity. Our results illustrate the practical need for robust corpuses of high-quality Windows PE binaries in training modern learning-based binary analyses. Assemblage can be downloaded from https://assemblage-dataset.net/.
1 Introduction
Binary analysis, the task of understanding and analyzing an executable binary program, is extraordinarily labor-intensive. Practitioners often spend hours to weeks manually analyzing a single program (Raff et al., 2020). The development of automated binary analysis tools has also proven to be extremely labor-intensive. Historically, such tools have relied on careful and judicious software engineering efforts to handle all corner cases, or used SAT solvers and similar data structures to resolve as many conditions as possible (Liu et al., 2023b). Given this high human effort, it is unsurprising that an increasing number of binary analysis methods use machine learning to help automate such efforts. This includes low-level tasks such as function disassembly (Yu et al., 2022), and high-level tasks like malware detection (Rokon et al., 2020) or variable name identification (Lacomis et al., 2019; Chen et al., 2022) which leverage these lower-level results.
However, significant issues in data availability have hampered the development of machine learning-based binary analysis for decades (Abt and Baier, 2014; Raff and Nicholas, 2020). Disassembly and decompilation require a mapping between the original true source or assembly code and the corresponding bytes of binary instructions, but it is non-trivial to collect and maintain such a mapping. Malware detection requires access to a collection of “benign” programs to train on, but can not be legally distributed due to copyright and license concerns. The situation is particularly bad for Windows PE binaries: while Windows is the most frequent target of malware, modern binary analysis systems train almost exclusively on Linux ELF binaries. Toolchain identification requires knowledge of the entire compilation process and of a variety of compiler settings that are simply not recorded in the original file. Binary function similarity requires having ground truth of functions that are semantically equivalent but written or compiled differently; most existing models rely on coding competition data that does not reflect the diversity of real-world code. The ad-hoc approaches taken by prior works to tackle these issues further complicate the reproducibility of these methods due to the infrastructure challenges in setting up an environment to build the same tools and perform complex data pre-processing (Marcelli et al., 2022; Raff, 2019, 2021; Gundersen et al., 2022). Moreover, authors often do not have the necessary licenses or do not make the code publicly available for others to extend the work.
We introduce Assemblage, a tool that makes significant progress towards overcoming all of these challenges. Assemblage is a framework for crawling code hosting platforms, downloading open-source projects, compiling them with multiple different compiler versions and settings, and recording the necessary information to reconstruct byte-level mappings from the output back to the source code. By using open-source code targets from GitHub, we obtain greater diversity in programs, resolve licensing/distribution issues, and provide detailed ground truth to support all of the aforementioned research areas. This alleviates the concerns with non-semantic preserving modifications (Fleshman et al., 2018) and the exorbitantly expensive cost and more limited diversity of binary re-writing (Wong et al., 2023; Boutsikas et al., 2021). Assemblage is designed to be extensible so that additional sources of code, compilers, and feature extractors can be added for further functionality. Assemblage datasets can be further distributed as a list of repositories and settings so that issues with mixed-license repositories (e.g., GPL and BSD) do not make a binary-based distribution unlicensable. Finally, we have designed and tested the system to ensure the reproducibility of datasets built by Assemblage.
Modern binary analysis systems use ever-complex models, with the recent switch to large-language models (LLMs) necessitating orders-of-magnitude more data given the rapidly increase in network size and number of parameters. In Section 2 we survey several of these efforts, observing that the largest system trains only 48k Linux ELF binaries. We designed Assemblage to produce large, diverse corpuses of binaries for training these state-of-the-art models at unprecedented scale. Our experiments running Assemblage over the past year on AWS has produced a large corpus of benign Windows PE (890k) and Linux ELF (428k) binaries. Assemblage clearly tracks source-to-binary provenance (down to the instruction level, in the case of Windows PE binaries) to provide a rich array of features and facilitate compliance with licensing and distribution. Assemblage’s design facilities a wide degree of customization and extension, allowing almost every aspect of its pipeline to be customized to suit the needs of modern, data-intensive binary analysis tools.
Our contributions include:
  • Assemblage, a cloud-based distributed system carefully engineered to automatically construct diverse corpuses of high-quality Windows PE binaries while ensuring compliance with licensing and terms of service.
  • An evaluation of Assemblage’s scalability and reproducibility, generating over 890k binaries over 20 weeks.
  • The Assemblage Windows and Linux datasets, consisting of 890k and 428k binaries respectively, with a total of 29 different combinations of CPU architecture, optimization level, compiler, compiler version, and other compiler flags.
  • Three case studies that demonstrate Assemblage’s application to (a) compiler provenance detection and (b) function similarity identification. Our case studies demonstrate the need for Assemblage’s rich features, particularly for corpuses of Windows PE binaries.
2 Related Work and Open Challenges
Prior Work: Crawling Open Source Repositories
Table 1: Binary corpuses—Assemblage vs. the literature
Dataset Binaries Functions Toolchain Dataset Format Public Assemblage-WinPE 890k 252M MSVC Binaries + PDB Assemblage-WinPE-Licensed 97k 41M MSVC Binaries + PDB Assemblage-LinuxELF 428k 316M GCC/Clang Binaries Assemblage-LinuxELF-Licensed 211k 186M GCC/Clang Binaries EMBER 1.1M undisclosed undisclosed Extracted features Features PassTell 552 149k GCC/Clang/ICC Extracted features DIRE 164k 3M GCC Binaries SPEC(XDA) 3.1k 9.75k GCC/ICC/MSVC Custom format GHCC (DIRTY) 75k  1M GCC Binaries VirusShare 67M undisclosed undisclosed Binaries BinaryCorp-26M(jTrans) 48k 25.8M GCC Binaries SOREL-20M 10M undisclosed undisclosed Binaries
Significant prior work has focused on crawling open-source repositories to build artifacts. For example, Sourcer (Linstead et al., 2007) crawled SourceForge to create a corpus of software for source-code similarity. More recent work targets GitHub due to its current dominance. This includes projects such as GHTorrent, which crawls and indexes GitHub (Gousios, 2013), along with other work focusing on improving crawling technology  (Barmpis et al., 2020). The open-source tool GHCC builds Linux-based GitHub repositories in batch mode and supports several other systems (such as DIRTY, XDA, and DIRE) (Lacomis and Hu, 2020).
Assemblage distinguishes itself from prior work by offering multi-node scalability, end-to-end reliability, an extensible interface, and deep integration with Windows toolchains, and a comparison is shown in Table 1. To the best of our knowledge, there is no preexisting work that provides a systematic and replicable solution to generate Windows binaries from a wide range of sources. In fact, much of the recent work applying machine learning techniques to binary analysis tasks (e.g. (Du et al., 2022; Wang et al., 2022)) relies on custom crawlers that cannot be extended to Windows PE binaries due to GUI-based toolchains.
The PE Binary Availability Crisis
The challenges of obtaining and distributing benign Windows PE executables continues to push modern work to rely upon manually-curated corpora of Linux binaries, or, alternatively, on paid services. For those who can afford it, VirusTotal has become the de-facto standard in malware classification due to its offering of benign binaries (Salem et al., 2021; Zhu et al., 2020b, a). For those who cannot, there are very few other options. While the distribution of malicious binaries is unfettered by licensing restrictions (e.g., VirusShare by Roberts (2011)), these restrictions can complicate distribution of benign corpuses of PE binaries, subsequently inhibiting reproducibility of machine-learning-based binary analysis tools. Additionally, even when data is available, it is often in preprocessed form (e.g., weights or features) that precludes its application to novel tasks. For example, while the popular EMBER data set includes 300k malicious and 300k benign Windows PE binaries, only EMBER’s features are public, not the binaries themselves (Anderson and Roth, 2018). While much work uses open-source Linux binaries (e.g., DIRE by Lacomis et al. (2019)), our case studies show that the utility of Linux-based training data for Windows-based tasks is unclear.
Consistent with our description of the landscape, there were many systems to which we were unable to apply Assemblage, due primarily to the lack of (a) available comparison data or code, or (b) support for Windows PE binaries. For example, work such as XDA makes data available only in preprocessed form (Pei et al., 2021). Similarly, tools such as PalmTree (a model for learning instruction embeddings) do not include implementations of the downstream tasks necessary to perform end-to-end comparison (Li et al., 2021b). Finally, almost all systems (such as Debin (He et al., 2018), Symlm (Jin et al., 2022), and DIRTY (Chen and Sun, 2022)) work only on Linux binaries, relying upon DWARF or ELF-specific details for feature extraction. We speculate that the development of machine-learning-based binary analysis systems is significantly hampered by the serious costs involved with obtaining large benign corpora of PE binaries.
Models Continue to Grow; Binary Datasets Must Keep Up
The most recent developments in binary analysis have been driven by the success of Large Language Models (LLMs) for natural language tasks. Encouraged by the impressive performance of models like BERT (Devlin et al., 2019a) and GPT-4 (OpenAI, 2023), researchers have adapted the methods of natural language processing to apply to the domain of binary executables. One branch of work has focused on building general assembly language models, which can be fine-tuned to perform a wide variety of tasks (Li et al., 2021b; Artuso et al., 2022; Koo et al., 2021; Zhu et al., 2023). Other work presents models with specific downstream tasks in mind, including binary code similarity detection (Wang et al., 2022; Ahn et al., 2022; Gu et al., 2023; Liu et al., 2023a), compiler provenance identification (He et al., 2022), vulnerability detection (Yu et al., 2021), and malware detection (Li et al., 2021a; Hamad et al., 2021; Wu et al., 2021).
While these LLM-inspired models have advanced the state of the art in many areas of binary analysis, they demand orders-of-magnitude more input data to avoid overfitting, magnifying the effect of PE binary availability crisis. LLMs derive their name partially from the fact that they are trained on massive amounts of data, and it is well known that LLMs can achieve significantly better performance and even unlock emergent abilities at scale (Brown et al., 2020; Wei et al., 2022). However, a surprising amount of existing work applies LLMs to relatively small input corpuses. We were able to find dataset breakdowns for 10 of 13 models mentioned in the previous paragraph; the average number of binaries used to compose the training set for these models is roughly over 13k (median 3.7k). Two models are trained on fewer than 200 binaries, with Liu et al. (2023a) being trained on only 48 binaries. Wang et al. (2022) and Zhu et al. (2023) are trained with the BinaryCorp dataset, which, at 48,130 (Linux) binaries, is more than twice the size of the next largest dataset used. We hope that the orders-of-magnitude greater amount of data made available by Assemblage (890k Windows PE and 428k Linux ELF binaries) will further improve the performance of these state-of-the-art tools.
Modern Efforts in Binary Analysis Ignore Windows
We found that nearly half of the transformer-based models discussed above are trained largely, or even primarily, on executables from standard Linux packages such as coreutils, binutils, diffutils, findutils, busybox, libcurl and openssl. While these packages are well-written and easy to access, compile, and parse, they do not align with the quality and characteristics typically found in the codes written by the majority of programmers. The over-reliance on common Linux packages for training exemplifies a major problem with these training datasets; out of the thirteen transformer-based models discussed, only one, Wu et al. (2021), is trained using Windows data. Moreover, the reliance on small collections of Linux binaries extends beyond transformer-inspired models (Marcelli et al., 2022; Kim et al., 2022; Al-Kaswan et al., 2023).
With the increasing use of binary analysis in security-relevant settings, the absence of publicly available corpora of benign Windows binaries is a critical concern. The difficulty of obtaining Windows PE data in academic literature is prevalent in related domains like malware detection that rely on small datasets (Joyce et al., 2022), effect sizes (Patel et al., 2023), and label noise (Joyce et al., 2021a, b, 2023), which generally depend on VirusShare and VirusTotal for data. An examination of files uploaded to VirusShare in 2023 indicates that 40.38% of the roughly 1.8 million malware samples are Windows binaries, whereas less than one percent are ELF binaries (Roberts, 2011). [^1] The other major file types are PDF, which accounts for roughly 25% of files, and Unknown, which accounts for 32% of files and captures files whose types could not be parsed. Files in the unknown category include data files, some scripts, ASC files, and source code. Given the prevalence of Windows executables, and the fact that Windows accounts for roughly seventy percent of global operating system market share (Statcounter, 2023), it is troubling that nearly all the recent literature on machine learning approaches to binary analysis uses Linux binaries as training data. This is especially worrying because, as we demonstrate in Sections 6.2 and 6.3, models trained on Linux binaries do not necessarily generalize to the Windows domain. Therefore, in order to produce useful machine learning models for binary analysis, it is necessary at a minimum to test, and likely necessary to train, these models on Windows data. We hope that will Assemblage fill the dataset gap to enable the application of modern efforts to Windows PE executables.
Comparing Assemblage and Preexisting Corpuses
Table 1 outlines a collection of binary corpuses from the literature, along with the largest corpus generated by Assemblage on top. The data set name is shown on the left, followed by the number of binaries, the number of functions (across all binaries), and the toolchains. The next column illustrates the dataset format: Assemblage provides a detailed source-to-binary mapping as part of the compilation process, and generates Program DataBase (PDB) files for use with reverse engineering tools such as IDA Pro. EMBER is a popular data set of Windows PE binaries, but the raw binaries are not public due to licensing restrictions; instead, EMBER provides only extracted features (Anderson and Roth, 2018). PassTell’s data set is also in the form of features, though its source is available and the basis of one of our case studies. While Windows PE binaries from XDA (generated by GHCC) are available, the total number of functions is relatively small (roughly 10k) (Pei et al., 2021); in more recent work using transformer learning, the XDA authors use Linux binaries gathered from open-source repositories to attain a sufficient amount of training data (Chen et al., 2022).
In sum, the vast bulk of work in modern binary analysis uses corpuses which are either (a) far too small for modern applications, or (b) consist exclusively of ELF binaries. Assemblage addresses the limitations of these existing datasets by providing large corpuses of both Linux and Windows binaries, complete with provenance and attribution to ensure compliance with licensing concerns and exposing the source-to-binary mapping.
3 Assemblage’s Design and Implementation
Figure 1: Assemblage architecture. A user downloads a corpus built via a worker pool on AWS, driven by the coordinator, which is configured via an interface accessed via secure shell.
We designed Assemblage as a cloud-based distributed system tailored to produce robust, reproducible, and high-quality binary corpuses. Figure 1 illustrates Assemblage’s principal components.
Distributed Architecture
Assemblage is designed to run as a distributed system on any cloud infrastructure, implementing a star topology system architecture. It consists of a central coordinator node, an adaptive worker pool, and a database interface. The system is designed to be resistant to component failure and endure high throughput. Assemblage provides rich APIs to enable wide flexibility in configuration, deployment, and extension.
Assemblage is driven by the coordinator node, which is responsible for configuration, task management, and database updating. Administrators communicate with Assemblage through a command-line interface provided by the coordinator. Then, a RabbitMQ message broker performs task management via a single-producer multi-consumer message queue. Finally, the coordinator dispatches tasks to specific channels of workers in the worker pool and, once a worker marks a task as complete, updates its database accordingly.
Assemblage’s worker nodes form an adaptive pool where each worker listens to one designated task channel. New workers can join the pool by registering with the coordinator and listening to a specific channel. Conversely, the coordinator can mark a worker as offline if it doesn’t send back a heartbeat signal within an allotted time period.
The modular design of Assemblage ensures that the unavailability of certain services does not result in a system-wide failure. The worker pool protects the system from cascading failures, so that the crash of one worker does not lead to a system-wide collapse. The message queue helps system communication remain stable as the worker pool scales. In practice, we have deployed up to 100 workers concurrently without issue.
Task Lifecycle
We will first describe our GitHub crawler, though Assemblage abstracts around input streams for source code, and we also include input sources for local (on-disc) repositories, Arch Linux’s AUR, and Microsoft’s vcpkg. The crawler begins by sending queries to GitHub’s APIs to discover new repositories. The coordinator then assembles results into build tasks, which instruct a worker to clone the repository and build it in a specific configuration. Workers listening to the channels will clone and build the binaries indicated in each task, and store the collected binaries in a designated storage location (typically an AWS S3 Bucket). Build workers are implemented by instantiating an abstract interface; we implemented builders for both Windows PE and Linux ELF binaries. On Windows, Assemblage uses the Microsoft Visual Studio toolchain, incorporating a set of automation scripts to overcome Visual Studio’s reliance on GUI-based user interfaces. On Linux, we implement workers for both GCC and Clang, using Docker to facilitate flexible and isolated deployment.
Tracking Build Provenance
Most prior work that built datasets of executables relied on disassembly tools to extract function information from binaries. For example, Wang et al. (2022), Ahn et al. (2022), and Marcelli et al. (2022) use IDA Pro to retrieve function information, while Du et al. (2022) uses objdump. Though significant engineering effort has gone into improving such disassembly tools over the years, they still do not work flawlessly, especially on Windows binaries. For example, IDA Pro can only infer function boundaries, but not names, on Windows PE binaries. Moreover, there is no guarantee as to the correctness of the recovered function information, especially when certain sections of information are stripped entirely(Wang et al., 2022). To address the shortcomings inherent in using tools such as IDA Pro, Assemblage is designed is carefully track build provenance, enabling the collection of ground truth compiler flags and settings.
Assemblage’s methods for recording build provenance are operating system dependent. For Linux binaries, we keep the compiled binary’s symbol table, unstripped, to preserve function information. This information is then added to the SQLite database corresponding to the dataset. For Windows binaries, in addition to collecting the binaries themselves, Assemblage also collects Program Database (PDB) files and other log files generated by the Microsoft Visual C++ compiler. These PDB files contain essential debug information that can be leveraged by binary analysis tools such as IDA Pro and Ghidra to associate source code with the bytes of compiled binaries and generate precise analysis outcomes. For the ease of generating training labels from PDB files, Assemblage processes PDB files utilizing Microsoft Debug Interface Access (DIA) SDK’s sample code “DIA2dump” (Microsoft, 2022a), then stores the files in exported Assemblage datasets. Additionally, the compilation metadata recorded in the PDB files is stored in the dataset’s SQLite database for portability.
By controlling the metadata passed to compiler during compilation, Assemblage ensures that the records in its SQLite table are accurate. To compare the utility of Assemblage’s metadata to that collected by previous methods, we passed a subset of Assemblage Windows binaries, containing a total of 23.7 million functions, to IDA Pro 8.2. IDA was unable to recover the names of nearly 10.7 million (45%) of those functions, instead returning a generic subroutine_location label (e.g. sub_10013040). The incidence of irretrievable function names in Assemblage’s metadata, by comparison, was $6\cdot 10^{-7}\%$. This demonstrates that Assemblage provides more complete and more accurate binary executable data than earlier dataset collection procedures.
Serializing “Recipes” for Reproducibility
The distribution of binary corpuses is complicated by the data size, licensing issues, and the need to accommodate application-specific features (e.g., control-flow graphs extracted via a commercial tool). To solve these issues, Assemblage supports distribution of datasets via “recipes,” compact Python bootstrap scripts that capture all information necessary to reconstruct a dataset. Given a recipe, users can replicate the state of a running Assemblage instance, as details about worker nodes and their associated instance types, build configurations, and repository URLs are all included. Recipes take advantage of Assemblage’s declarative API to enable wide flexibility in system operation. For example, via a small change in the recipe, Assemblage can be switched to either continuously crawl for new repositories, replicate an existing set of repositories, or build a set of repositories from a local disk. This allows sharing a smaller “recipe” over a larger data dump and a backup mechanism in the event of data corruption or hardware failure.
4 The Assemblage Datasets
890k Windows-Github PE Dataset
We constructed a dataset of 890K Windows PE binaries as follows: starting from 4 million crawled GitHub repositories dating from 2008 to 2023, we discarded repositories that had fewer than 10 files, were smaller than 50KB in size, lacked a README.md file containing the string “Visual Studio”, or lacked a solution file supporting Microsoft Visual Studio tool chains. We were left with 179,548 repositories, from which Assemblage was then able to build and store binaries using our Microsoft Visual Studio toolchains. Build configurations varied in toolchain version (Visual Studio 2015, 2017, 2019, and 2022), optimization level (0d, 01, 02, 0x), CPU architecture, and Debug/Release mode. In sum, we ended up with 847,856 executable (.exe) files and 43,507 dynamically-linked library (.dll) files, along with PDB files for each binary. In total, we accumulated 2TB worth of data, of which the executables (.exe and .dll files) took 200GB. To conform with ethical data distribution practices, we recorded license information from crawled repositories when available. Table 2 shows the distribution of licenses present among the GitHub repositories in our corpus; consistent with our observation of many small repositories, 158,624 of the 179,548 repositories have no license.
Table 2: Top licenses in GitHub datasets.
License Windows-GitHub Linux-GitHub
MIT 7258 6762
GPL 3078 2845
Apache 1168 859
BSD 313 609
CC 94 50
BSL 22 49
Figure 2: Binary size distribution of Windows on GitHub, vcpkg and Linux on GitHub data.
We measured our dataset along a variety of dimensions to demonstrate its quality and diversity. Figure 2 shows a histogram of sizes of binaries from our (a) Windows binaries from GitHub, (b) vcpkg (c) Linux binaries from GitHub corpuses. Our Windows-GitHub binaries were the smallest overall, with an average size of 143KB. Manual inspection revealed many student projects, basic utilities, and starter codes among our GitHub dataset. However, in absolute terms, our Windows-GitHub dataset still contains a substantial number of large binaries, including 26,549 binaries over 0.5MB.
Table 3: Most common error codes from Assemblage Windows dataset
Error Code Ratio Memo
MSB4126 47% Invalid specified solution configuration
C1083 18% Missing source file
MSB8036 8% Specified Windows SDK version was not found
MSB3202 3% Missing project file
MSB4025 3% The imported project file loading error
We also present a breakdown of our Windows-Github dataset by compiler and optimization level (Table 4). The number of binaries at each compiler version and optimization is not equal, due to unevenly distributed computing power among Assemblage worker nodes, changes to the availability of repositories, restricted running time, and the limited availability of some configurations in solution files. To further investigate the varying number of binaries under each build configuration, we recorded the error message encountered during build failures, which we summarize in Table 3. Over the whole dataset, the most common causes of build failure are related to compiler infrastructure, such as invalid solution configurations, Windows SDK errors due to upgrading, invalid or missing project files, and other compilation errors. Although we tried to mitigate the errors caused by missing files by installing a wide range of SDK libraries, complete coverage was infeasible, and thus certain library versions were absent from our worker environment. We also did our best to minimize invalid configuration errors by minimally modifying variables in vcproj files and using devenv from MSVC to upgrade solution files when appropriate.
Table 4: Configuration distribution on three datasets.
OS Source Compiler Opt. Count
Windows PE GitHub MSVC-v140 O1 60126
O2 165866
Ox 100831
MSVC-v141 O1 6113
O2 78203
Od 86841
Ox 7470
MSVC-v142 O1 87497
O2 50055
Od 72613
MSVC-v143 O1 80243
Od 57011
Ox 37302
vcpkg MSVC-v120 O1 952
O2 945
Od 956
Ox 926
MSVC-v142 O1 2160
O2 2142
Od 2187
Ox 2145
MSVC-v143 O1 3081
O2 3078
Od 3074
Ox 3083
Linux ELF GitHub GCC-11.4.0 Od 50338
O1 49860
O3 50995
Oz 22030
Clang-14.0.0 Od 50338
O1 63188
O2 65934
O3 69082
Moreover, we manually inspected the log files associated with certain build configurations that had particularly low success rates: v141-O1-Debug-x64, v141-Ox-Debug-x64, and v140-O1-Release-x86. We present the results in Table 5. We found that 86% of v141-O1-Debug-x64 and 66% of v141-Ox-Debug-x64 failures produce error code MSB4126, which indicates that the solution configuration is invalid. By comparison, the dominant error code for v140-O1-Release-x86 is C1083, which indicates that a certain file is missing or wrong.
Figure 3: Function distribution of each binary.
We also inspected the function information corresponding to our Windows-Github PE dataset. We found that out of 252M functions, there are 14M unique function names, and 138M unique function hashes. Assemblage is also able to store source code for 17M functions in its SQL database. We also observed significant variation in the number of functions per binary, which we illustrate in Figure 3. Finally, we examined the relationship between a binary’s size and the functions it contains. Counter-intuitively, the number of functions in a binary is not correlated with the size of that binary, as these variables have a Pearson correlation coefficient of 0.22. Furthermore, the size of functions within a binary has a 0.49 correlation coefficient to the binary size, but no correlation to the number of functions in the binary. We unexpectedly conclude that binary size doesn’t have a strong correlation with either the number of functions nor the average function size in a binary. Moreover, the existence of function level information provides insights into the behavior patterns of specific compilers. For example, in 2017 Visual Studio implemented stack-based buffer protection by inserting security cookies, and we can locate more than 744k __security_init_cookie functions with 151k unique hashes. These function level information provided by Assemblage presents valuable data for not only machine learning research, but also a potential of application to a wide range of research areas like security and binary analysis.
Table 5: Error codes for specific build configuration
Flags MSB4126 C1083 MSB8036 MSB4025 Others
MSVC-v140 O1-Release-x86 25% 34% 0% 1% 40%
O2-Debug-x64 82% 5% 0% 0% 13%
O2-Release-x86 25% 34% 0% 1% 40%
Ox-Release-x64 21% 36% 0% 0% 43%
MSVC-v141 O1-Debug-x64 86% 3% 0% 4% 6%
O2-Release-x86 30% 24% 0% 6% 40%
Od-Release-x86 37% 21% 0% 6% 35%
Ox-Debug-x64 66% 10% 0% 6% 18%
MSVC-v142 O1-Release-x64 20% 23% 30% 1% 26%
O2-Release-x86 25% 22% 28% 1% 25%
Od-Debug-x64 69% 7% 12% 0% 12%
MSVC-v143 O1-Debug-x86 25% 18% 33% 6% 18%
Od-Debug-x64 84% 4% 2% 4% 6%
Ox-Release-x86 18% 36% 16% 2% 28%
25k Windows vcpkg Dataset
Upon examining our Windows-Github dataset, we observed that only 4.7% of the corpus consisted of library files. To address this deficiency and provide coverage of Windows’ statically-linked and dynamically-linked libraries, we adapted Assemblage to build binaries from vcpkg (Microsoft, 2022b), a dataset of 2,185 actively-developed, well-maintained, and widely-used packages. vcpkg consists of many commonly used C++ libraries, including graphics and UI-related libraries. Additionally, every project hosted on vcpkg is licensed, which makes this dataset easy to redistribute.
Our vcpkg worker takes advantage of the native commands of vcpkg, and we modified the package manager’s original build script to pass custom variables and flags to its compiler. In this way, we harvested a diverse Windows PE library corpus from vcpkg, comprising 24905 binaries, along with their pdb files, from 2078 different projects. Only 110 of these projects are also present in our Windows-Github dataset, demonstrating the added diversity that is present in our vcpkg dataset. Similarly to the Windows-Github dataset, vcpkg binaries were built using 4 different versions of the Microsoft Visual Studio toolchain, as well as with 4 optimization levels (Od, O1, O2, O3). Finally, as illustrated in Figure 2, the size of the binaries in our vcpkg dataset is larger than the size of those in our Windows-GitHub dataset. The vcpkg binaries were 2124 KB on average, compared to 143KB for Windows-Github binaries, with a small number (roughly 500) of vcpkg binaries measuring more than 10 MB.
Proof-of-Concept Linux Dataset
Motivated by the scarcity of corpuses of benign Windows PE binaries, we have focused our efforts on building such executables. However, to ensure Assemblage’s API was sufficiently expressive, we also added several Linux builders. Our prototype Linux crawler traverses GitHub and searches for Makefiles to identify potential Linux ELF projects. Among 4 million crawled repositories, Assemblage recognized 367,207 repositories that include a Makefile. We created 8 build configurations, four using GCC (-O0, -O1, -O3, and -OZ) and four using Clang (-O0, -O1, -O2, and -O3).
Similar to what we did when building our Windows PE database, we deployed Assemblage to AWS with the same coordinator, but with workers on different EC2 instances. In sum, Assemblage generated 428,884 Linux binaries from 30,856 GitHub repositories. The histogram of the file size distribution of our Linux ELF dataset is shown in Figure 2, the composition of compiler and optimization level of the dataset is presented in Table 4, and the overall dataset size is recorded in Table 1. Interestingly, we observe a much higher frequency of licensed repositories in our Linux-Github dataset; nearly half of our Linux-Github binaries have licenses, compared to only 11% of our Windows-Github binaries.
Generally, our Linux builder failed more often than our Windows PE builder (a success rate of 7% compared to the 30% on Windows), but produced larger files in the projects that it did build. Though it is natural to be concerned about build failures during the dataset creation process, these errors are not introduced by Assemblage, as we simply use regular expression to substitute the compiler flags stored in the Make files, then call make to start the building. Indeed, a manual inspection confirmed that most of the build errors resulted from errors in the repository’s source code or Make file.
In addition to the GitHub crawler, we also built a proof-of-concept crawler and builder that leverages the Arch User Repository (AUR) for source code. Users can set the configuration by providing custom makepkg.conf for each project. While the AUR repositories are well maintained by the community and open-sourced, due to the rate limits and ethical concerns, we did not deploy builders to generate a large scale dataset from AUR. However, our proof-of-concept crawler demonstrates that Assemblage can be configured to work with a variety of package managers in order to harvest binaries.
Assemblage’s SQLite Database
A typical Assemblage dataset consists of both (a) an SQLite database of metadata and (b) an archive of binaries (either as a .tar or manually via the Amazon S3 interface). The SQL database provides a logical specification of a collection of binary files, currently .exes and .dlls. The user queries the SQLite database to perform application-specific feature extraction (e.g., identifying pairs of sibling functions for triplet learning), and often feeds the binaries into a machine learning framework such as TensorFlow, PyTorch, or Keras.
Figure 4: Dataset Database Schema
The SQLite database consists of four primary tables: binaries, functions, rvas, and lines (illustrated Figure 4). The binaries table retains a reference of each binary’s configuration used during compilation, e.g. platform, build mode, compiler version, and optimization, and assigns each binary an unique identifier for further reference. We also register additional information about each executable including the file name, the location of the project on GitHub, the project’s license, and the size of the binary in this table. Using the binaries table, researchers can curate subsets of the dataset to meet their training needs, e.g. by selecting only binaries of a certain size or with a certain license.
To provide more knowledge on functions and minimize the need to distribute pdb files (since, as we saw earlier, pdb files can take up 90+of dataset space), the functions, rvas and lines tables utilize the unique identifier for each binary stored in binaries table to present the relationship between function level information and binaries. The functions tables offers a high-level overview of functions, presenting a hash of the function’s bytes and the function name, whereas the rvas and lines tables provide lower-level function details. Specifically, the rvas table documents the relation between each function and the its relative virtual address when the binary is loaded into memory. On the other hand, the lines table illustrates the mapping of each source code line to the precise byte address located in the compiled binary. The dataset also provides the relation between binaries and their corresponding pdb files in a pdbs table, in case researchers want to make use of those files for further binary analysis.
Figure 4 shows the information that is available about a single binary within an Assemblage dataset. As an example, we query the database for a binary called butler.exe, which has the unique identifier 184456. In the binaries table, we get basic facts about the executable, including the compiler version (Visual Studio 2017), build mode (Release), optimization flag (O2), and associated license (MIT). In the functions table, we can see that this binary has 568 functions; in particular, we focus on one function called tinyxml2::StrPair::SetStr, uniquely identified by the tag 51629995. We use this unique identifier to query the lines and rvas tables for further information about this function. In the lines table, we can see the source code corresponding to this function, such as the line if (*(pu + 0) == TIXML_UTF_LEAD_0. From the rvas table, we learn that this function will be loaded into the relative addresses from Ox6460 to Ox64AA. Thus, this example demonstrates how all the information needed to do binary analysis with corpuses of Assemblage executables is available in the accompanying Assemblage SQLite database.
Limitations
It is natural to be concerned about the presence of malware in any dataset of binary executables. Assemblage accumulates binaries by compiling source code from GitHub, which explicitly prohibits malware on its platform and allows users to report suspected malware to its community discussion board. Currently, Assemblage does not independently verify the benignness of the binaries it collects, though we are looking to add malware scanning, following the model pioneered by EMBER (Anderson and Roth, 2018), in the next iteration of the tool.
5 Evaluating Assemblage on AWS
We deployed Assemblage on AWS for easy scalability. In this section, we detail our experience running Assemblage and generating datasets on this platform.
Cost and Performance of Assemblage over 20 Weeks
We ran Assemblage for over a year on AWS; here, we report on a systematic study of its cost over a 20 week period in 2023. The recipe chosen for this study, which aimed to build a diverse Windows PE binary corpus, used our crawler to identify C++ repositories (queried via language in GitHub’s API) created between January 2009 and July 2023. We gathered roughly 4 million suitable repositories, then attempted to build each of the repositories in 17 distinct configurations via Assemblage. While many repositories failed to build (e.g., Linux repositories), we ultimately built 791,564 Windows PE binaries from 161,796 repositories.
Figure 5: 20 Week Deployment Cost (USD)
Figure  5 shows the per-week cost to run Assemblage over our 20-week period; cost details were gathered from AWS’s “Cost Explorer” panel. The total cost for generating and warehousing our corpus of 890k binaries was \$12,285. Our cluster was initialized with 16 AWS T3.large Windows instances (\$0.11 USD/hour) as builder nodes, one T3.xlarge Linux instance (\$0.17 USD/hour) running the coordinator node and crawler, and one R6g.large RDS database instance to host a MySQL database of metadata.
Costs varied over the twenty-week period as we administered the cluster flexibly to best accommodate our varying needs. After two weeks of stable builds, we gradually doubled the number of worker nodes we deployed. During week 5, we realized the system was bottlenecked both due to differing success rates among our builders and due to GitHub-imposed limitations on cloning; thus, we scaled the system down slightly. By week 10 we had exhausted all available input repositories from GitHub. In week 14, we re-queued repositories that had previously failed to clone; this restored high-speed binary collection, prompting us to expand the cluster by adding more worker nodes.
Overall, the blue bars representing the instance running cost dominates the overall cost, and the S3 cost (purple portion) remained relatively low throughout most weeks, with occasional spikes as a result of dataset exporting. The cost of the database shown in green remains relatively constant for majority of the time, as there was no change in either size or quantity of database instances.
Evaluating Recipe Reproducibility
Repositories on GitHub are constantly updating with new commits and changes in branches. This revolving process can cause the binaries produced by compiling latest commit of source code to be different from those compiled using previous versions. Our reproducibility study measured the impact of the dynamic nature of GitHub on Assemblage’s performance. We measured Assemblage’s reproducibility by reconstructing a 10k Windows PE binary dataset from one recipe three times. The collected binaries from each round were compared to evaluate consistency and reliability.
We deployed 4 T3.large builder nodes with 2 CPU cores (Intel Xeon Platinum 8000 series) and 8 GB of RAM, 1 T3.xlarge coordinator node with 4 cores (Intel Xeon Platinum 8000 series) and 16 GB RAM, and a R6g.large RDS database. Each evaluation on generated dataset involved running the builds simultaneously on each node, with 6 building processes for each build configuration. In consideration of bandwidth, cloning time, and relatively low possibility of the repositories being pushed with new commits in our experiment time span, we only do a shallow git clone, that is to only retrieve the latest commit without history git information. Assemblage stores git commit hash and last update time of each repository in recipe, and users can use this information to clone the whole repository and revert back to the last recorded commit, to ensure the compiled source code doesn’t change over new commits. To select our 10k binaries, we first used Assemblage to randomly crawl and build repositories from GitHub until 10,000 binaries were successfully built; then we created a recipe at the moment and reconstructed the dataset from the same recipe three separate times.
Table 6: Reproducibility Study (Three Rounds)
Round 1 2 3
Total Task Count 10,000 10,000 10,000
Successful Task Count 10,000 10,000 10,000
Zip File Count 10,000 10,000 10,000
Binary Count 11,177 11,177 11,177
Running Time (hrs) 10.4 10.9 10.4
The results of our three-round data reproduction testing are presented in Table 6 based on the database, storage service and binary dataset. In three replication experiments, the total 10,000 tasks are all successfully built and yielded the same amount of binaries. We also inspected the binary file and source code and ensured that each binary are built from same source code and mostly identical (due to the nature of compilers, even building on same source code will yield slightly different binaries). Overall, our experiments confirmed the reproducibility and determinism of Assemblage as a dataset generation tool.
6 Assemblage for ML: Three Case Studies
In this section, we demonstrate the utility of Assemblage for a variety machine learning tasks. We use Assemblage Windows datasets to evaluate three widely-regarded machine learning approaches to binary analysis problems: a LightGBM model for compiler provenance identification (Du et al., 2022), a graph neural network (GNN) for binary function similarity detection (Li et al., 2019), and a BERT-based transformer network for understanding binary files, fine-tuned on the task of binary function similarity detection (Wang et al., 2022). Our experience highlights the challenges of applying current machine learning models to Windows data, and the commonness of these challenges across model architectures and tasks. Specifically, we demonstrate that models developed on Linux data do not perform well when applied to Windows data, signifying a lack of generalizability, and we note the prevalence of Linux-specific tool chains in the literature. Our findings reveal the importance of Assemblage datasets, both as a tool for evaluating existing techniques and as a data source for training the next generation of machine learning models.
6.1 Case Study: Learning Compiler Provenance
Compiler provenance inference is a machine learning task for determining the compiler configuration used to build a binary, including compiler version, build flags, and compiler passes. We trained a state-of-the-art provenance learning model PassTell (Du et al., 2022) on a combination of data, mixing the authors’ original Linux dataset with Windows PE binaries obtained from Assemblage. This allowed us to evaluate PassTell’s performance on a more diverse corpus, demonstrating the utility of Assemblage in assessing binary analysis models.
We first used Assemblage to build a dataset of Windows binaries that complemented the Linux binaries studied in the original PassTell paper. We attempted to build each of the 15 source projects selected by Du et al. (2022), but due to differences between the Linux and Windows platforms and their respective compiler infrastructures, some projects could not be rebuilt. In total, we collected 127 binaries, which we proceeded to preprocess via the same protocol as the original paper. Though the PassTell authors used objdump, a linear disassembler, to extract function bytes, we found higher success rates with IDA Pro, augmented with the PDB files collected by Assemblage, to accurately handle the significant number (5%) of discontiguous functions present in Windows binaries. Following disassembly, we masked operands and literals in the same manner as Du et al. (2022), and balanced the training data by compiler toolchain, version, and optimization level.
Figure 6: PassTell’s Confusion Matrix on the dataset mixing the author’s training data (top left) with Assemblage’s Windows PE binaries (bottom right).
We conducted two experiments using PassTell, one with mixed-platform data, and one with only Windows data. For our mixed-platform experiment, we combined the Clang and GCC binaries from the original PassTell dataset with our dataset of Windows binaries, which were built with the Microsoft Visual C++ Compiler (MSVC We then re-trained PassTell on our mixed dataset on an Ubuntu 22.04 server equipped with an AMD EPYC 7713P@1.9Ghz 64-Core Processor and 512GB of RAM. The confusion matrix in Figure 6, which depicts the results of this experiment, displays an ambiguous diagonal pattern, reflecting the model’s recognition capacity for compiler optimizations. Individual compilers are distinguishable as square areas, indicating the model’s ability to discern among the three compilers. Note that the outline of the square formed by the MSVC compiler is much more prominent than those of the other two compilers, yet the diagonal of the confusion matrix is least pronounced in that section. This indicates that PassTell performs much worse at compiler version and optimization level identification on the Windows portion of the dataset than it does on its Linux counterpart. Additionally, when comparing our Linux results with the authors’, we observed significantly lower prediction accuracy for Clang binaries, where precision drops to around 60%. This could be because the original paper’s analysis does not cover cross-function optimizations such as the function inlining on Clang data.
Figure 7: PassTell’s Confusion Matrix on the Assemblage’s Windows PE Binaries.
To better determine how much the difference in operating systems affected PassTell’s accuracy on the mixed dataset, we conducted a second experiment using solely our Windows data. The result is shown as the confusion matrix in Figure 7. Although there is a recognizable diagonal line, indicating that the model has the ability to tell the difference between various compiler versions and settings to some degree, the overall accuracy peaks at 0.58, lower than the accuracy on GCC and Clang that the authors claim in the paper. We conclude that the approach proposed by the PassTell authors could not effectively capture the difference between versions and optimization levels from MSVC.
The observation of PassTell’s performance on Windows data indicates the limits of training models on small binary datasets lacking diversity. While the approach put forth by Du et al. (2022) seemed promising when applied to a small Linux dataset, extending the work to Windows data from Assemblage revealed serious shortcomings. This underscores the utility of Assemblage as a tool for machine learning researchers in the binary analysis space.
6.2 Case Study: Binary Function Similarity
Next, we used Assemblage to evaluate how well a graph neural network (GNN) trained on Linux data generalizes to Windows binaries. We conducted these experiments on the GNN developed by Li et al. (2019) for binary function similarity detection, as implemented and open-sourced by Marcelli et al. (2022). [^2] Code from https://github.com/Cisco-Talos/binary_function_similarity/. We adhered closely to the procedures outlined in the Marcelli et al. (2022) README, making use of their scripts and notebooks to create a dataset of Assemblage functions preprocessed according to the specifications of the model from Li et al. (2019). We made only one change to their data processing pipeline — instead of removing singleton functions, we allowed them to stay in the dataset and be part of negative pairs, as they can still provide a training signal to the model this way. By following the same data processing protocols as Marcelli et al. (2022), we ensured a fair comparison with their results.
Table 7: The criteria for positive pairs for each task. An X indicates that the designated variable must differ in the pair. A $\ast$ indicates that the designated variable may differ. If a cell is unmarked, then its corresponding variable does not differ.
Arch Bit Compiler Version Opt
arch X
bit X
comp X X
ver X
opt X
XA X X
XC X X X
XC + XB X X X X
XM $\ast$ $\ast$ $\ast$ $\ast$ $\ast$
We started with a dataset of $219,265$ binaries collected by Assemblage. After removing functions with fewer than five basic blocks and discarding duplicate functions (identified by comparing a hash of their opcodes), we were left with a final dataset of $1.78$ million functions. From here, we created ground truth positive and negative pairs of functions on which to evaluate our model. Negative pairs consist of two different functions, whereas positive pairs consist of functions originating from the same source code, but differing in one or more of the following: architecture, bitness, compiler, compiler version, and optimization. Marcelli et al. (2022) identifies nine different evaluation tasks, separated by the nature of the positive pair (see Table 7). Because Assemblage data is currently restricted to a single architecture and compiler, we were only able to create evaluation pairs for four of the nine tasks: bit, ver, opt, and XM. From our $1.78$ million functions, we were able to create 6 positive pairs for task “bit”, $22,434$ positive pairs for task “opt”, $21,031$ positive pairs for task “ver”, and $298,880$ positive pairs for task “XM”, which we subsampled down to $50,000$ pairs. For each task, we created the same number of negative pairs as positive pairs. Because we only had 6 positive pairs for “bit”, we chose not to evaluate this task, as there wasn’t sufficient data to present a meaningful result.
Table 8: AUC scores evaluating binary function similarity model generalization between the Linux and Windows domains.
Train Windows Test Linux Train Linux Test Windows Train Linux Test Linux
arch 0.50 0.97
bit 0.70 0.99
comp 0.63 0.80
opt 0.72 0.62 0.88
ver 0.82 0.64 0.98
XA 0.48 0.86
XC 0.63 0.86
XC+XB 0.61 0.87
XM 0.54 0.61 0.87
We report AUC scores for three sets of experiments. In the first experiment, we used the $1.78$ million Windows functions in our Assemblage dataset to train the GNN model published in Marcelli et al. (2022) from scratch, then tested on their original Linux test dataset. To facilitate comparison, we kept the same training settings used by authors, training the embedding model in pair mode for ten epochs with opcode features. In the second experiment, we tested the Linux pre-trained GNN model from Marcelli et al. (2022) on the Windows function pairs created from the Assemblage dataset. Finally, the results from running the Marcelli et al. (2022) Linux-trained GNN against the Linux test set are included as a baseline. Results from all experiments are presented in Table 8.
The results in this case study reinforce our conclusions from our study of compiler provenance identification. As the scores in Table 8 make clear, the binary function similarity detection GNN from Li et al. (2019) is unable to transfer knowledge between the Linux and Windows contexts. Due to the ubiquity of Windows software, many downstream applications of binary analysis tools, including function similarity detection and provenance identification tools like the ones tested here, will be to Windows binaries. Therefore, it is essential that these tools perform well when applied to Windows files. As the experiments in this section demonstrate, this necessitates a large amount of labeled, high-quality Windows data for these tools to be developed on. Assemblage addresses this need by providing the first robust technique for curating high-quality Windows data.
6.3 Case Study: Transformer-Based Learning
In addition to traditional techniques for machine learning, we also evaluated our Assemblage dataset on some more recent transformer-based approaches for binary code similarity detection. jTrans (Wang et al., 2022) is a recently-proposed transformer method for binary code similarity. Based on the BERT paradigm (Devlin et al., 2019b), jTrans uses masked language modeling and a novel task, jump target prediction, in its pre-training phase. This is followed by fine-tuning on labeled function data using triplet loss.
The jTrans authors released their data processing scripts, and we were able to use their code to process our Assemblage data, with one exception. To extract function byte addresses, the jTrans script exploits features specific to Linux binaries. To accommodate Windows PE binaries into processing, where this technique wouldn’t apply, we instead used the function information contained in Assemblage’s SQLite database to retrieve function byte addresses. Other than that, we precisely followed the script released by the jTrans authors to create a training and validation set.
Table 9: Reproducing jTrans on our Windows and Linux datasets
Completed Epochs 1 20 50
Fine-tune & Evaluate on Windows 0.165 0.519 0.52
Fine-tune & Evaluate on Linux 0.829 0.829 0.83
Fine-tune & Evaluate on Linux (as reported in jTrans paper) 0.797 N/A 0.97
The jTrans authors did not release their pre-training code, so we were only able to conduct fine-tuning experiments on Windows and Linux data. We randomly selected 72000 Windows PE binaries and 23000 Linux ELF binaries from Assemblage’s dataset, processed them, and separated them into training and validation sets for Windows and Linux, respectively. For both the Windows and Linux datasets, using the hyperparameter settings recommended by the jTrans authors, we ran fine-tuning for 50 epochs, at which point the MRR over the training set became stable. Then we used the withheld validation set to calculate the MRR score on 4 optimization levels; the result is shown in Table 9.
Table 9 shows a notable difference between the reported MRR scores on Linux (0.83) and Windows (0.52) data, even with fine-tuning. Pre-training dominates the model’s learning in our Linux experiments, and fine-tuning on Linux data doesn’t contribute much to the model’s performance, increasing accuracy on Linux test data by only .001 percent after fifty epochs. On the other hand, fine-tuning on Windows data raised the model’s MRR score from 0.17 to 0.52 on Windows test data. Because the authors have not released their code for pre-training, we could not conduct any further investigation into the impact of the pre-training step. However, it is significant that jTrans and jTrans-zero (jTrans without fine-tuning) both perform worse on the larger, more diverse Assemblage data than they do on their original training corpuses. This demonstrates the value that Assemblage brings to researchers, who can leverage its datasets to conduct experiments on more challenging and more realistic data.
7 Conclusion
Binary analysis is a rapidly-accelerating field whose growth is placing increasing demands on the availability of diverse, high-quality binary corpuses. Unfortunately, binary analysis has a data availability crisis: licensing challenges, file size, and lack of standardization in data set distribution continue to obfuscate the comparison of state-of-the-art work.
To help build the next generation of robust datasets of binaries, we introduced Assemblage, a distributed system that crawls, configures, builds, and distributes richly-labeled binary corpuses in a variety of configurations. To our knowledge, Assemblage is the largest-ever effort to build publicly-available corpuses of Windows PE binaries. Assemblage encapsulates a wide variety of corpus construction methodologies via a “recipe,” which contains all information necessary to precisely reproduce a binary corpus. We demonstrate Assemblage’s robustness and scalability by studying its performance over a 20-week period in 2023, producing 890k Windows PE binaries, totaling over 258M functions. Following that time, we built a Linux-based builder using GCC and Clang, and built 428k Linux ELF binaries across eight configurations.
Given the prevalence of the Windows operating system, we believe that a primary use case of machine learning tools for binary analysis will be on Windows binaries. The vast bulk of ML algorithms in this space are currently trained and evaluated solely on Linux binaries; our case studies illustrate the challenges of generalizing state-of-the-art models trained on Linux ELF binaries to work on PE binaries. For example, we observe in our measurements of Passtell that models that perform well on small GCC-based binaries may fare much worse on larger Windows PE binaries (see Figure 6); our measurements of binary function similarity paint a similar picture. We conclude that Assemblage provides the first large scale dataset of Windows binaries. It can easily be leveraged to train models for a variety of binary analysis tasks, and fills a critical gap in the field and enables new research directions. Along these lines, we are currently working to use our datasets to perform language modeling, leveraging the rich source-to-binary mapping available in Assemblage to apply transformer-based language models to Windows binaries for the first time.
We present Assemblage as a remedy to the dataset gap in binary analysis, and hope that it will facilitate reproducible infrastructure for evaluating the next generation of binary analysis tools, built on increasingly-complex models.